A home was bought with a 20% down payment of R177175,00.The balance was financed for 20 years at 9,4% interest per annum,compounded semi-annually.Find the size of the half-yearly payments for the loan


1
Expert's answer
2021-10-12T02:11:11-0400

Let the cost of the home be x,

Therefore, 20% of x= R177175

X= R885875, which is the total payment.

Balance= R885875-R177175

= R708700.

Determination of half years installment payment using the formula below;

"P=A[\frac{1-{(1+\frac{r}{m})^{-n}}}{\frac{r}{m}}]"

Where P= balance or future payments= R708700, A= Installment payment half years, r= interest rate per annum= 9.4% or 0.094, n= number of years= 20years and m= compounding periods (half yearly=2.)


"R708700=A[\frac{1-{(1+\frac{0.094}{2})^{-20}}}{\frac{0.094}{2}}]"

"A= R9203.796587"


The size of the half-yearly payment for the loan is R9203.796587